How much does the gas cylinder last? Volume=14l, Pressure=200 bar and flux=6l/min.
1
Expert's answer
2020-12-10T09:40:13-0500

"PV = constant \\

P_1V_1 = P_2V_2 \\

200 \times 14 = 1 \times V_2 \\

V_2 = 2800 \;L \\

Flux=6 \;l/min \\

Time \; T = \frac{V_2}{Flux}

= \frac{2800}{6} \\

= 466.6 \;min = 7.77 \;hr"

Answer: 466.6 min or 7.77 hr